OpenStudy (anonymous):

Derivative of f(x) = sqrt(x) - 5/sqrt(x) + 5x^7/3+3-2/X^3.

zepdrix (zepdrix):

the square roots giving you a little trouble? :o

OpenStudy (anonymous):

\[f(x)=\sqrt{x} - \frac{ 5 }{ \sqrt{x} } + 5x ^{\frac{ 7 }{ 3 }} + 3 + \frac{ 2 }{ x^{3} }\] Math Syntax.

OpenStudy (anonymous):

So far I understand: \[\sqrt{x} = \frac{ 1 }{ 2 }x ^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

-2/x^3 < Correction.

zepdrix (zepdrix):

hmmm looks good, so how about the second term? \[-5x^\frac{ -1 }{ 2 }\] The important thing to keep in mind with the power rule is: When you're decreasing negative powers, they're actually increasing in the negative direction.

zepdrix (zepdrix):

\[(d/dx)x^{1/2}=(1/2)x^{-1/2}\] like you said the first time.

OpenStudy (anonymous):

Oh I mean for the last bit after the stand alone 3 should be minues not plus. The math syntax post had that typo.

zepdrix (zepdrix):

oh oh oh XD

zepdrix (zepdrix):

lemme give you the last term, and maybe it'll help ya figure out the second term :O \[(d/dx)\frac{ -2 }{ x^3 }=(d/dx)-2x^{-3}=(-2)(-3)x^{-4}\] make sense? :o

OpenStudy (anonymous):

Taking the reciprocals?

zepdrix (zepdrix):

\[\frac{ 1 }{ x^3 }=x^{-3}\] is that the part that's confusing you? :o Yah you bring the x into the numerator, and apply a negative exponent to it.

OpenStudy (anonymous):

Is it then -5x^2/3 ?

OpenStudy (anonymous):

-2/3 That is.

OpenStudy (anonymous):

-3/2 I mean gah.

OpenStudy (anonymous):

Considering a -1/2 - 1 = -3/2.

zepdrix (zepdrix):

\[(d/dx)\frac{ -5 }{ \sqrt x }=(d/dx)-5x^{-1/2}=(-5)(-1/2)x^{-3/2}\] Yah your exponent looks correct :)

OpenStudy (anonymous):

I'm going to follow up with my answer soon.

OpenStudy (anonymous):

Ok, so I went for a meal break, Would this be correct:\[\frac{ 1 }{ 2 }x ^{\frac{ -1 }{ 2 }} - (-5)(\frac{ 1 }{ 2 })x ^{\frac{ -3 }{ 2 }} + 5(\frac{ 7 }{ 3 })x ^{\frac{ 4 }{ 3 }}-(-2)(-3)x ^{-4}\] Which then after changing signs: \[\frac{ 1 }{ 2 }x ^{\frac{ -1 }{ 2 }} + 5(\frac{ 1 }{ 2 })x ^{\frac{ -3 }{ 2 }} + 5(\frac{ 7 }{ 3 })x ^{\frac{ 4 }{ 3 }} +2(-3)x ^{-4}\] ?

zepdrix (zepdrix):

And I assume taking the derivative of the +3 gave you 0? Yah looks correct :) good job.

OpenStudy (anonymous):

Yep, constants = 0 so +/- 0 just eliminates the 3. Thank you so much. I feel this forum is really great for help. I am going to contribute to others hopefully as much as I learn from others. :D